A customer service call center handles 5100 calls in an 8-hour day. If the average length of time to handle each call is 8 minutes, and nobody is expected to be out sick or on vacation, at least how many agents working the same 8-hour shift would you need to handle all calls? How many agents do you need if the average length of time to handle each call is 6 minutes?
Answer by rfer(16322)

8*60=480
480/8=60
5100/60=85 agents
------------
480/6=80
5100/80=63.75=64 agents